OpenStudy (anonymous):

write in simplest form 2cosx/sin2x

OpenStudy (minecrafter64):

??????

OpenStudy (mathmale):

Time to haul out our table of trig identities. If you'd find a suitable identity for sin 2x, you'll probably see right away how you could reduce 2cos x / sin 2x.

OpenStudy (anonymous):

use the formula for sin2x: \[\sin 2x = 2\sin x \cos x\] in the denominator. Then, cancel the cosines to get 1/(2sinx) = (csc x)/2.

OpenStudy (anonymous):

\[\frac{ 2cosx }{ \sin2x }=\frac{ 2cosx }{ 2sinxcosx }=\frac{ 1 }{ sinx }=\sin^{-1}x\]

OpenStudy (anonymous):

=cscx

OpenStudy (anonymous):

oops. forgot about the 2 in the numerator...:)

OpenStudy (anonymous):

thank you

OpenStudy (mathmale):

\[\frac{ 2cosx }{ \sin2x }=\frac{ 2cosx }{ 2sinxcosx }=\frac{ 1 }{ sinx }=\sin^{-1}x\] is almost, but not quite, right. Up to 1/sin x, great. But

OpenStudy (mathmale):

\[\frac{ 1 }{ \sin x }\neq \sin ^{-1}x.\] Rather, \[\frac{ 1 }{ \sin x }= \csc x\]

OpenStudy (mathmale):

and \[\sin ^{-1}x \] is a function in its own right: the INVERSE SINE FUNCTION.

OpenStudy (anonymous):

indeed @mathmale